6 g of a hydrated salt on heating gives 4.4 g of of anhydrous salt.if the molecular mass of an hydrous salt is 150, the water of hydration is

Grade:11

3 Answers

Shashank
5 years ago
Let the water of hydration be x.
Now mass of water=6 - 4.4=1.6g.
Moles of salt in LHS= 6/150
Therefore,moles of water in =6x/150
Mass of water=(6x/150)18
And in reaction we have find that mass of water is 1.6g(...in 2nd line..)
Therefore,(6x/150)18=1.6g
                x=2.2222..
So,water of hydration is 2
Riya Kalam
4 years ago
 M xH2O ----------->      M     +    xH2O              
(150+8x/6)                (150/4.4)                    ,
therefore, on cross multipling ,  [ (150 + 18x)4.4  =  150 × 6 ]            
                                                               ie,, x=3
Sarasi T
3 years ago
mass of hydrated salt =6 g
mass of anhydrous salt =4.4 g
mass of water formed = 1.6 g
molecular mass of anhydrous salt =150 g
M .xH2O on heating gives M +xH2O
no. of moles =wt/molecular mass
molecular mass of hydrated salt = 150+18x
no. of moles of M.xH2O(Hydrated salt)= 6/(150+18x)
no. of moles of anhydrous salt =4.4/150
therefore; on equating the two moles relation : 6/(150+18 x) =4.4/150
                                                                            900=660 +79.2 x
                                                                                x =3.03
                                                                                 x = 3